What NaCl concentration results when 259ml of a 0.810 M NaCl solution is mixed with 562 ml of a 0.280 M NaCl solution?
1
Expert's answer
2020-03-02T06:56:05-0500

We will use the formula

"MV=M_1V_1+M_2V_2"

"M(259+562)=0.810\times259+0.280\times562 \newline
M(821)=209.8+157.4=367.20"

"M=0.45"